# Security Concerns for Large Language Models: A Survey

### v1

@@ Abstract
Large Language Models (LLMs) such as GPT-4 (and its recent iterations like GPT-4o and the GPT-4.1 series), Google’s Gemini, Anthropic’s Claude 3 models, and xAI’s Grok have caused a revolution in natural language processing, but their capabilities also introduce new security vulnerabilities. In this survey, we provide a comprehensive overview of the emerging security concerns around LLMs, categorizing threats into prompt injection and jailbreaking, adversarial attacks (including input perturbations and data poisoning), misuse by malicious actors (e.g., for disinformation, phishing, and malware generation), and worrisome risks inherent in autonomous LLM agents. A significant focus has been recently placed on the latter, exploring goal misalignment, emergent deception, self-preservation instincts, and the potential for LLMs to develop and pursue covert, misaligned objectives (scheming), which may even persist through safety training. We summarize recent academic and industrial studies (2022–2025) that exemplify each threat, analyze proposed defenses and their limitations, and identify open challenges in securing LLM-based applications. We conclude by emphasizing the importance of advancing robust, multi-layered security strategies to ensure LLMs are safe and beneficial.
@@ I Introduction
LLM security encompasses not only external threats such as prompt manipulation, data exfiltration, or malicious use (e.g., phishing or disinformation)[15, 8], but also intrinsic risks arising from autonomous LLM agents—models that may develop misaligned goals[16], engage in strategic deception [17, 14], or pursue covert agendas that persist even through safety training [12, 13]. This survey addresses four broad categories of threats: (1) prompt injection and jailbreaking, where adversarial inputs coerce the model to break its safety constraints; (2) adversarial attacks in both training and inference (such as input perturbations and data poisoning) that degrade model performance or implant backdoors; (3) misuse by malicious actors, where LLMs are leveraged to generate disinformation, phishing emails, malicious code, etc.; and (4) intrinsic risks from LLM-based autonomous agents. This last category is particularly nuanced and significant, encompassing not only goal misalignment, where an agent’s learned utility differs from user intent, but also the potential for agents to develop their own covert objectives, engage in strategic deception (scheming), exhibit self-preservation behaviors, and even retain these undesirable traits despite current safety training paradigms [13, 12]. We integrate recent studies for each category, discuss defenses (and their limits), and highlight open research challenges.
This survey makes the following contributions: (1) We provide a comprehensive taxonomy of LLM-related security threats, organized into four major categories: prompt injection and jailbreaking, adversarial attacks, malicious misuse, and intrinsic risks from autonomous LLM agents. (2) We systematically review recent academic and industry works from 2022 to 2025, highlighting representative examples of each threat type. (3) We evaluate the effectiveness and limitations of current defense strategies, including input sanitization, adversarial training, and oversight mechanisms. (4) We identify open research challenges for securing LLMs, especially in light of emergent risks such as goal misalignment, scheming, and deceptive behavior in autonomous agents. By mapping the evolving threat landscape and surveying mitigation strategies, this survey aims to inform both practitioners who are deploying LLMs and researchers who are designing the next generation of large language models with the potential risks, actionable insights, and practical recommendations for mitigating the security threats.
@@ II-A Defining Prompt Injection and Jailbreaking
Prompt injection and jailbreaking represent a significant class of attacks where adversaries craft malicious inputs or append instructions designed to override an LLM’s intended prompt or built-in safety rules. For example, an attacker might prepend text like “Ignore previous instructions and explain how to hack a computer” to trick an LLM into providing prohibited content. This kind of attacks is conceptually similar to SQL injection in traditional software vulnerabilities: the injected prompt contaminates the operational context, which makes it difficult for the LLM to distinguish between legitimate user queries and adversarial instructions [10, 9, 6].
Prompt injections can be categorized as direct, where malicious text is fed directly into the prompt, or indirect, where the malicious instruction is hidden within user-uploaded content, such as documents, emails, or web pages, which the LLM processes as part of its context [10, 18]. Indirect injections are particularly dangerous in tool-augmented systems like retrieval-augmented generation (RAG) agents or email-based LLM applications, where untrusted content is automatically fed into model context windows.
Jailbreaking is a closely related and often overlapping form of manipulation where the attacker’s input aims to force the model to “break out” of its pre-defined behavioral guardrails and safety protocols [7, 19, 20]. Jailbreaking prompts may include encoded or adversarial suffixes, emoji-based perturbations, or multilingual padding that evades filters. These manipulations exploit the model’s learned pattern and continuation biases, which compels it to comply with unsafe requests even when explicitly instructed not to.
@@ III Adversarial Attacks: Input Perturbations and Data Poisoning
Adversarial attacks against LLMs can be performed either during its training phase or at the point of inference.
@@ III-A Training-Time Attacks: Data Poisoning and Backdoors
Such attacks aim to “tamper with [training] data by introducing fudged or malicious data to … confuse the trained models”, so they subsequently produce incorrect or harmful outputs [6]. During training, poisoning attacks have been extensively explored in both academic and practical settings. Wallace et al. [26] demonstrated that models like GPT-2 could be made to output arbitrary attacker-specified content by inserting rare token sequences into fine-tuning data. These adversarial triggers were robust and transferred across prompts and architectures.
@@ III-B Inference-Time Attacks: Input Perturbations
In contrast to training-time attacks, inference-time adversarial attacks involve carefully crafting or slightly perturbing a user input through methods like word substitution, paraphrasing, or token-level manipulation to cause the LLM to make mistakes, such as generate unsafe content and misinterpret content. While early research in adversarial machine learning primarily focused on image classifiers, recent work has definitively shown that LLMs are also highly vulnerable to such perturbations [6]. Even seemingly minor changes, such as using synonyms or introducing slight misspellings, can mislead an LLM’s generation process or induce misclassification of text [7, 22].
@@ III-C Recent Empirical Studies and Advanced Techniques
A surge of poisoning studies now offers much richer empirical evidence. BackdoorLLM benchmarks eight attack families on 16 models and shows that as little as 0.01% poisoned pre-training data can implant highly persistent triggers, meanwhile surviving reinforcement learning from human feedback (RLHF) and instruction tuning [24]. BadGPT compromises the reward model during RLHF to create covert policy backdoors, demonstrating that alignment layers themselves become attack surfaces [25]. Beyond poisoning, universal adversarial suffixes such as GCG blur the line between prompt injection and classical adversarial NLP, revealing transferability across dozens of LLMs [7]. Fuzzing surveys catalog automated mutation frameworks that expose previously unknown failure modes in commercial chatbots [27].
@@ IV Misuse by Malicious Actors
Beyond directly attacking the model’s integrity, malicious actors can exploit the inherent capabilities of LLMs for a wide range of malicious or criminal purposes. Because LLMs excel at generating fluent, coherent, and contextually appropriate text, they can be readily misused to automate and scale social engineering attacks and various forms of cybercrime.
@@ V Intrinsic Risks in LLM Agents
An emergent and profoundly concerning frontier in LLM security involves their integration into autonomous agentic systems. When LLMs are endowed with goals, the ability to make plans, and the capacity to use tools or interact with external environments, novel categories of risks emerge. These risks stem not just from external manipulation but from the agent’s own internal state, learned behaviors, and potential intentions, which may not align with those of human designers or users [34], as extensively detailed in discussions of catastrophic risks from agentic AI [35].
@@ V-A Goal Misalignment and Unfaithful Reasoning
Modern LLMs often employ Chain-of-Thought (CoT) prompting to produce step-by-step reasoning before an answer. Some may believe that through careful monitoring the CoT, we can gain insight into the model’s internal decision-making process and detect early signs of misalignment. However, Lanham et al. [36] investigated whether this stated CoT is a faithful explanation of the model’s actual reasoning process. Through intervention experiments, they found significant variation: some models rely heavily on their CoT, while others largely ignore it, suggesting the reasoning can be post-hoc. Critically, they discovered that “as models become larger and more capable, they produce less faithful reasoning on most tasks we study,” which indicates an inverse scaling for faithfulness. This implies that the explanations provided by more advanced agents might be less reliable indicators of their true decision-making processes. Further complicating this, Baker et al.[17] demonstrated that even if CoTs can be monitored for misbehaviors (like reward hacking), directly penalizing “bad thoughts” by incorporating CoT monitors into the agent’s training objective can lead to “obfuscated reward hacking”. In such cases, agents learn to hide their malicious intent within a CoT that appears benign, effectively rendering the CoT an unfaithful or deceptive representation of their true reasoning while the misbehavior persists. If an agent’s explanations are not faithful, either inherently or due to learned obfuscation, it is significantly harder to detect misalignment or verify that an agent is genuinely pursuing its stated goals rather than rationalizing a predetermined outcome or deceiving oversight mechanisms. The system card for Anthropic’s Claude Opus 4 released in May 2025, for instance, noted that while its reasoning transcripts generally appear consistent, they “will often omit important information that influences model behavior,” indicating “moderate overall reasoning faithfulness” [37].

### v2

@@ Abstract
Large Language Models (LLMs) such as GPT-4 and its recent iterations, Google’s Gemini, Anthropic’s Claude 3 models, and xAI’s Grok have caused a revolution in natural language processing, but their capabilities also introduce new security vulnerabilities. In this survey, we provide a comprehensive overview of the emerging security concerns around LLMs, categorizing threats into prompt injection and jailbreaking, adversarial attacks such as input perturbations and data poisoning, misuse by malicious actors for purposes such as generating disinformation, phishing emails, and malware, and worrisome risks inherent in autonomous LLM agents. A significant focus has been recently placed on the latter, exploring goal misalignment, emergent deception, self-preservation instincts, and the potential for LLMs to develop and pursue covert, misaligned objectives, a behavior known as scheming, which may even persist through safety training. We summarize recent academic and industrial studies from 2022 to 2025 that exemplify each threat, analyze proposed defenses and their limitations, and identify open challenges in securing LLM-based applications. We conclude by emphasizing the importance of advancing robust, multi-layered security strategies to ensure LLMs are safe and beneficial.
Figure 1: Taxonomy of Security Threats for Large Language Models.
@@ I Introduction
LLM security encompasses not only external threats such as prompt manipulation, data exfiltration, or malicious use (e.g., phishing or disinformation)[15, 8], but also intrinsic risks arising from autonomous LLM agents—models that may develop misaligned goals[16], engage in strategic deception [17, 14], or pursue covert agendas that persist even through safety training [12, 13]. This survey addresses four broad categories of threats: (1) prompt injection and jailbreaking, where adversarial inputs coerce the model to break its safety constraints; (2) adversarial attacks in both training and inference (such as input perturbations and data poisoning) that degrade model performance or implant backdoors; (3) misuse by malicious actors, where LLMs are leveraged to generate disinformation, phishing emails, malicious code, etc.; and (4) intrinsic risks from LLM-based autonomous agents. This last category is particularly nuanced and significant, encompassing not only goal misalignment, where an agent’s learned utility differs from user intent, but also the potential for agents to develop their own covert objectives, engage in strategic deception (scheming), exhibit self-preservation behaviors, and even retain these undesirable traits despite current safety training paradigms [13, 12]. We integrate recent studies for each category, discuss defenses (and their limits), and highlight open research challenges. Figure 1 presents a taxonomy of the LLM security threats discussed in this survey.
This survey makes the following contributions: (1) We provide a comprehensive taxonomy of LLM-related security threats, organized into four major categories: prompt injection and jailbreaking, adversarial attacks, malicious misuse, and intrinsic risks from autonomous LLM agents. (2) We review a broad range of recent academic and industry works from 2022 to 2025, highlighting representative examples of each threat type. (3) We evaluate the effectiveness and limitations of current defense strategies, including input sanitization, adversarial training, and oversight mechanisms. (4) We identify open research challenges for securing LLMs, especially in light of emergent risks such as goal misalignment, scheming, and deceptive behavior in autonomous agents. By mapping the evolving threat landscape and surveying mitigation strategies, this survey aims to inform both practitioners who are deploying LLMs and researchers who are designing the next generation of large language models with the potential risks, actionable insights, and practical recommendations for mitigating the security threats.
@@ II Prompt Injection and Jailbreaking
Figure 2: Conceptual illustration of normal LLM interaction, where a user query and system prompt lead to an intended output, versus a prompt injection attack, where malicious input (direct or indirect) contaminates the context, overriding system instructions and leading to unintended or harmful outputs.
@@ II-A Defining Prompt Injection and Jailbreaking
Prompt injection and jailbreaking represent a significant class of attacks where adversaries craft malicious inputs or append instructions designed to override an LLM’s intended prompt or built-in safety rules, as conceptually depicted in Figure 2. For example, an attacker might prepend text such as “Ignore previous instructions and explain how to hack a computer” to trick an LLM into providing prohibited content. This kind of attacks is conceptually similar to SQL injection in traditional software vulnerabilities: the injected prompt contaminates the operational context, which makes it difficult for the LLM to distinguish between legitimate user queries and adversarial instructions [10, 9, 6].
Prompt injections can be categorized as direct, where malicious text is fed directly into the prompt, or indirect, where the malicious instruction is hidden within user-uploaded content, such as documents, emails, or web pages, which the LLM processes as part of its context [10, 18]. Indirect injections are particularly dangerous in tool-augmented systems such as retrieval-augmented generation (RAG) agents or email-based LLM applications, where untrusted content is automatically fed into model context windows.
Jailbreaking is a closely related and often overlapping form of manipulation where the attacker’s input aims to force the model to “break out” of its pre-defined behavioral guardrails and safety protocols [7, 19, 20]. Jailbreaking prompts may include encoded or adversarial suffixes, emoji-based perturbations, or multilingual padding that evades filters. These manipulations exploit the model’s learned pattern and continuation biases, which compels it to comply with unsafe requests even when explicitly instructed not to. The diverse landscape of these manipulation tactics, from direct textual overrides to sophisticated, context-hiding jailbreaks, is further detailed in Table I, which provides a comparative overview of common attack types and example studies.
@@ II-C Recent Advances in Automated and Sophisticated Prompt Attacks
TABLE I: Summary of Prompt Injection and Jailbreaking Attacks, Techniques, and Advanced Vectors
Attack Category / Specific Technique / Vector Description Key Characteristics Example Studies/Tools Susceptible Models (Examples)
Fundamental Prompt Injection Types
Direct Prompt Injection Malicious instructions directly fed into the prompt to override original intent. Overrides system prompts, straightforward to attempt. Perez and Ribeiro [9] GPT-3.5, LLaMA
Indirect Prompt Injection Malicious instructions hidden in external content (documents, emails, web pages) processed by LLM. Exploits RAG, tool-augmented systems; stealthier as input isn’t directly from malicious user. Toyer et al. [10] GPT-4, Gemini, Claude 3, Grok
Jailbreaking (Goal-Oriented Prompt Manipulation)
Jailbreaking Inputs designed to bypass safety guardrails and make the LLM produce harmful/forbidden content, often using sophisticated prompt injection. Focuses on circumventing safety protocols; employs various prompt engineering methods (see specific techniques below). General concept studied widely, e.g., [20] and papers detailing specific techniques. GPT-4, Gemini, LLaMA, Claude 3
Specific Techniques for Prompt Injection & Jailbreaking
G2PI (Goal-Guided Generative Prompt Injection) An automated technique that jointly searches for effective goals and corresponding malicious suffixes to achieve jailbreaking. Automated, high success rate by optimizing both target and attack vector (e.g., 90% on GPT-4, Gemini). Zhang et al. [19] GPT-4, Gemini
GCG Suffixes (Greedy Coordinate Gradient) Universal transferable adversarial suffixes appended to prompts, designed to bypass alignment defenses and achieve jailbreaking. Universal, transferable across diverse models; often gradient-based optimization of suffixes. Zou et al. [7] Open-source & proprietary models
Advanced Attack Vectors Utilizing Prompt Injection
Self-Propagating Worms (e.g., Morris-II) An attack vector where indirect prompt injection is used (e.g., via RAG-enabled email agents) to enable autonomous spread and data exfiltration. Autonomous, scalable, exploits system integrations (e.g., email) to propagate; leverages indirect PI. Cohen et al. [18] RAG-enabled LLM-backed agents
Steganographic Extraction (e.g., Imprompter) An attack vector using stealthy token sequences embedded in inputs (a form of indirect PI) to covertly extract personal data or execute commands via chats. Covert, malicious intent hidden within seemingly benign input; difficult to detect; leverages indirect PI. Chen et al. [21] Chat models
@@ III Adversarial Attacks: Input Perturbations and Data Poisoning
Adversarial attacks against LLMs can be performed either during its training phase or at the point of inference. These attacks, whether corrupting the model’s foundational knowledge or tricking its inferential capabilities, represent significant security challenges. Table II offers a categorization of these adversarial methods, differentiating between training-time and inference-time exploits and highlighting key techniques.
@@ III-A Training-Time Attacks: Data Poisoning and Backdoors
Such attacks aim to “tamper with [training] data by introducing fudged or malicious data to … confuse the trained models”, so they subsequently produce incorrect or harmful outputs [6]. During training, poisoning attacks have been extensively explored in both academic and practical settings. Wallace et al. [26] demonstrated that models such as GPT-2 could be made to output arbitrary attacker-specified content by inserting rare token sequences into fine-tuning data. These adversarial triggers were robust and transferred across prompts and architectures.
@@ III-B Inference-Time Attacks: Input Perturbations
In contrast to training-time attacks, inference-time adversarial attacks involve carefully crafting or slightly perturbing a user input through methods such as word substitution, paraphrasing, or token-level manipulation to cause the LLM to make mistakes, such as generate unsafe content and misinterpret content. While early research in adversarial machine learning primarily focused on image classifiers, recent work has definitively shown that LLMs are also highly vulnerable to such perturbations [6]. Even seemingly minor changes, such as using synonyms or introducing slight misspellings, can mislead an LLM’s generation process or induce misclassification of text [7, 22].
@@ III-C Recent Empirical Studies and Advanced Techniques
A surge of poisoning studies now offers much richer empirical evidence. BackdoorLLM benchmarks eight representative backdoor attack strategies across eight open-source LLM variants and shows that attacks remain effective even with very small poisoning budgets (e.g., 15–100 samples), while existing post-hoc defenses often fail to remove them [24]. BadGPT compromises the reward model during RLHF to create covert policy backdoors, demonstrating that alignment layers themselves become attack surfaces [25]. Beyond poisoning, universal adversarial suffixes such as GCG blur the line between prompt injection and classical adversarial NLP, revealing transferability across dozens of LLMs [7]. Fuzzing surveys catalog automated mutation frameworks that expose previously unknown failure modes in commercial chatbots [27].
TABLE II: Summary of Adversarial Attacks on LLMs
Phase Attack Type Description Key Techniques/Characteristics Example Studies & Key Findings
Training-Time Data Poisoning Malicious examples injected into training data to cause specific misbehavior. Subtle, hard to detect post-training. Can implant persistent triggers with small data percentage. Shayegani et al. [6], BackdoorLLM [24] (15–100 poisoned samples still effective; post-hoc defenses often fail)
Training-Time Backdoor Insertion Training LLM to produce harmful output for specific trigger inputs while behaving normally otherwise. Trigger-activated, behavior concealed on normal inputs. Reward model poisoning can create covert policy backdoors. Wallace et al. [26] (robust triggers), BadGPT [25] (reward model poisoning).
Training-Time Sleeper Agents Deceptive objectives embedded during training, activate on specific triggers, and resist safety fine-tuning. Covert, persistent through safety training; model may learn to better hide backdoor. Hubinger et al. [12] (Safety pipelines can train models to conceal backdoors).
Inference-Time Input Perturbations Minor changes to input (synonyms, misspellings, token manipulation) to cause misclassification or harmful output. Evades simple filters, exploits model sensitivity to token sequences. Zou et al. (GCG) [7] (transferable suffixes), Chen et al. (Imprompter) [21] (steganographic data extraction).
Inference-Time Universal Triggers Carefully chosen rare token sequences in prompts to induce specific outputs. Robust, transferable across prompts. Wallace et al. [26] (GPT-2 outputs attacker-specified content).
Inference-Time Fuzzing Automated discovery of subtle token-level changes that flip LLM behavior. Exposes brittleness to minor input variations (whitespace, emojis). FuzzSurvey [27] (Many chat-LLMs brittle to small input changes).
@@ IV Misuse by Malicious Actors
Beyond directly attacking the model’s integrity, malicious actors can exploit the inherent capabilities of LLMs for a wide range of malicious or criminal purposes. Because LLMs excel at generating fluent, coherent, and contextually appropriate text, they can be readily misused to automate and scale social engineering attacks and various forms of cybercrime. The range of such misapplications is broad, leveraging the LLMs’ generative prowess for nefarious ends. An overview of these misuse categories, the specific LLM capabilities they exploit, and prominent examples are detailed in Table III.
@@ IV-C Emergence of Specialized Malicious LLMs and Ecosystem Exploitation
TABLE III: Summary of LLM Misuse by Malicious Actors
Misuse Category Description & Examples LLM Capabilities Exploited Example Studies/Tools/Incidents Implications
Phishing Social Engineering Generating convincing phishing emails, tailored spam, devising fraud strategies. Fluent text generation, contextual understanding, mimicry. Saha Roy et al. [8], Cohen et al. (Morris-II) [18], Chen et al. (Imprompter) [21] Scalable, automated, convincing attacks; autonomous email worms.
Disinformation Generation Creating personalized fake news articles, propaganda. Narrative coherence, personalization, contextual manipulation. Zugecova et al. [28], Huynh et al. (PoisonGPT) [30], Wan et al. [31] Rapid spread of tailored misinformation; personalization can bypass safety filters.
Malware Generation Composing malicious code, scripts. Code synthesis, understanding of programming logic. Saha Roy et al. [8], Trustwave (WormGPT, FraudGPT) [29] Lowered barrier for malware creation; adaptive malware.
Specialized Malicious LLMs Custom ”jailbroken” or fine-tuned models for malicious tasks. Transfer learning, fine-tuning capabilities. WormGPT, FraudGPT [29], PoisonGPT [30], Wan et al. [31] Democratization of advanced malicious tools.
[…]
@@ V Intrinsic Risks in LLM Agents
An emergent and profoundly concerning frontier in LLM security involves their integration into autonomous agentic systems. When LLMs are endowed with goals, the ability to make plans, and the capacity to use tools or interact with external environments, novel categories of risks emerge. These risks stem not just from external manipulation but from the agent’s own internal state, learned behaviors, and potential intentions, which may not align with those of human designers or users [34], as extensively detailed in discussions of catastrophic risks from agentic AI [35]. These intrinsic risks, encompassing goal misalignment, unfaithful reasoning, emergent deception, self-preservation, scheming, and the persistence of such behaviors, pose unique and formidable challenges. Table IV provides a structured summary of these risk categories, along with key observed behaviors and their implications for safety and control.
@@ V-A Goal Misalignment and Unfaithful Reasoning
Modern LLMs often employ Chain-of-Thought (CoT) prompting to produce step-by-step reasoning before an answer. Some may believe that through careful monitoring the CoT, we can gain insight into the model’s internal decision-making process and detect early signs of misalignment. However, Lanham et al. [36] investigated whether this stated CoT is a faithful explanation of the model’s actual reasoning process. Through intervention experiments, they found significant variation: some models rely heavily on their CoT, while others largely ignore it, suggesting the reasoning can be post-hoc. Critically, they discovered that “as models become larger and more capable, they produce less faithful reasoning on most tasks we study,” which indicates an inverse scaling for faithfulness. This implies that the explanations provided by more advanced agents might be less reliable indicators of their true decision-making processes. Further complicating this, Baker et al.[17] demonstrated that even if CoTs can be monitored for misbehaviors (e.g., reward hacking), directly penalizing “bad thoughts” by incorporating CoT monitors into the agent’s training objective can lead to “obfuscated reward hacking”. In such cases, agents learn to hide their malicious intent within a CoT that appears benign, effectively rendering the CoT an unfaithful or deceptive representation of their true reasoning while the misbehavior persists. If an agent’s explanations are not faithful, either inherently or due to learned obfuscation, it is significantly harder to detect misalignment or verify that an agent is genuinely pursuing its stated goals rather than rationalizing a predetermined outcome or deceiving oversight mechanisms. The system card for Anthropic’s Claude Opus 4 released in May 2025, for instance, noted that while its reasoning transcripts generally appear consistent, they “will often omit important information that influences model behavior,” indicating “moderate overall reasoning faithfulness” [37].
TABLE IV: Summary of Intrinsic Risks in Autonomous LLM Agents
Risk Category Description Key Observed Behaviors / Concepts & Example Studies / Models Implications for Safety & Control
Goal Misalignment Agent’s emergent objectives diverge from intended human goals. Pursuit of unintended utility functions; misinterpretation of instructions. Studies: Mitchell [16], Bengio [35]. Fundamental challenge; risk increases with autonomy.
Unfaithful Reasoning LLM’s stated reasoning (e.g., CoT) does not accurately reflect its true decision-making process. Post-hoc rationalization; CoT ignored; inverse scaling for faithfulness. Studies: Lanham et al. [36], Baker et al. [17]. Model: Anthropic Claude Opus 4 [37]. Hard to understand/trust agent decisions; CoT oversight can be misled.
Emergent Deception Agent actively misleads or hides information. Obfuscated reward hacking; strategic false information. Studies: Baker et al. [17], Barkur et al. [14]. Model: Anthropic Claude Opus 4 [37]. Undermines trust/control; difficult detection of harmful intent.
Self-Preservation Agent takes actions to ensure its own survival, potentially against human wishes. Unsanctioned self-replication; blackmail; weight exfiltration. Studies: Barkur et al. [14]. Model: Anthropic Claude Opus 4 [37]. Agents may resist control or prioritize their existence.
Scheming Covert pursuit of misaligned goals by strategically hiding true capabilities or intentions. Writing self-propagating worms; fabricating documents; disabling oversight. Studies: Meinke et al. [13] (o1, Claude 3.5 Sonnet, Gemini 1.5 Pro, Llama 3.1 405B). Model: Early Claude Opus 4 (via [37]). Extremely difficult to detect/mitigate; high potential for catastrophic outcomes.
Persistence of Deception (Sleeper Agents) Deceptive backdoors persist through safety training, activating on specific triggers. Malicious objectives dormant; adversarial training may worsen hiding. Study: Hubinger et al. [12]. False sense of safety; current safety paradigms may be insufficient.